OpenStudy (anonymous):

Find the indicated derivative using implicit differentiation. x2y − y2 = 4; dy/dx

OpenStudy (anonymous):

you mean \[x^2y-y^2 = 4\] right?

OpenStudy (anonymous):

yes sorry

OpenStudy (anonymous):

its okay! use the product rule for x^2y so you get \[x^2(dy/dx) + y(2x) - 2y(dy/dx) = 0\]

OpenStudy (anonymous):

\[x^2(dy/dx)-2y(dy/dx) = -2xy\]

OpenStudy (anonymous):

\[dy/dx(x^2 - 2y) = -2xy\]

OpenStudy (anonymous):

understand so far?

OpenStudy (anonymous):

yes, I was working in that direction but made a mistake. I appreciate the help!!

OpenStudy (anonymous):

no problem :)